Practicing the Object Modeling Technology in a

Functional Programming Framework

Computation Structures Group Memo 379 March, 1996

Jack B. Dennis and Handong Wu

This memo is an extended version of a paper prepared for OOPSLA-96, Object-Oriented Programming Systems, Languages and Applications. Handong Wu is a visitor to the Laboratory for Computer Science on leave from Ericsson Telecom, Sweden.

The research reported in this docment was performed, in part, using facilities of the Laboratory for Computer Science of the Massachusetts Institute of Technology. Funding for the Laboratory is provided in part by the Advanced Research Projects Agency of the Department of Defense under the Office of Naval Research contract N00014-92-J-1310.